# Random pairing MLE for estimation of item parameters in Rasch model

### v1

@@ 4 Analysis
In this section, we present the main steps in the analysis to obtain theoretical results in the previous section. Section 4.1 gives a complete argument on the reduction to the BTL model we alluded in Section 2.2, Section 4.2 provides the analysis of the $\ell_{\infty}$ error, and Section 4.3 provides the analysis of both the $\ell_{2}$ error and the asymptotic distribution as they share similar components in the analysis.
4.1 Reduction to Bradley-Terry-Luce model
@@
A key idea in $\mathsf{RP\text{-}MLE}$ is the randomized pairing in Steps 1 and 2 of Algorithm 1. It compiles the user-item responses $\bm{X}$ to item-item comparisons $\bm{Y}$. In this section, we make a detailed argument that $\bm{Y}$ follows the Bradley-Terry-Luce model with a non-uniform sampling scheme.
@@
Recall that $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}$ and $Y_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}<X_{tj}\}$. The following fact provides the distribution of $Y_{ij}^{t}$ conditional on $L_{ij}^{t}=1$. We defer its proof to Section A.
@@
    Randomly split the $n_{t}$ problems taken by tester $t$ into $\lfloor n_{t}/2\rfloor$ pairs of problems. (Step 1(a) of Algorithm 1)
@@
    For all $(t,i,j)$, items $i$ and $j$ are compared by user $t$ if $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}=1$.
@@
4.2 Analysis of $\ell_{\infty}$ error
We have seen that analyzing $\mathsf{RP\text{-}MLE}$ under the Rasch model can be reduced to analyzing the MLE under the BTL model. This reduction allows us to invoke the general theory of MLE in the BTL model established in the recent work [YCOM24].
@@ 4.2 Analysis of $\ell_{\infty}$ error
The following lemma adapts Theorem 3 of the recent work [YCOM24] to our setting.
Lemma 1 (Theorem 3 in [YCOM24]).
@@
$\frac{np}{4\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{L})\leq\lambda_{1}(\bm{L}_{L})\leq 3np$ (13)
and
@@
4.2.1 Proof of Theorem 1
Now we are ready to prove Theorem 1. We focus on analyzing $\mathsf{RP\text{-}MLE}$, as the analysis of $\mathsf{MRP\text{-}MLE}$ follows immediately from the union bound of the different data splitting and the triangular inequality:
@@ 4.2.1 Proof of Theorem 1
4.3 Analysis for refined $\ell_{2}$ error and uncertainty quantification
To make the main text concise, we sketch the key ideas that lead to both Theorems 2 and 3.
We start with some necessary notation. For convenience, for any $(i,j)$, renumber the set of comparisons $\{Y_{ji}^{t}\}_{t:L_{ij}^{t}=1}$ as $\{Y_{ji}^{(l)}\}_{l=1,\ldots,L_{ij}^{t}}$ in an arbitrary order. Let $\epsilon_{ij}^{(l)}\coloneqq Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})$ where $\sigma(x)=e^{x}/(1+e^{x})$ is the sigmoid function. Let $\bm{B}\coloneqq\left[\cdots,\sqrt{z_{ij}}(\bm{e}_{i}-\bm{e}_{j}),\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y}}\in\mathbb{R}^{m\times L_{\mathrm{total}}}$ (repeat $L_{ij}$ times for edge $(i,j)$), where $L_{\mathrm{total}}\coloneqq\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}L_{ij}$ is the total number of observed effective comparisons in $\mathcal{G}_{Y}$. Define
$\widehat{\bm{\epsilon}}\coloneqq\left[\cdots,\epsilon_{ij}^{(l)}/\sqrt{z_{ij}},\cdots\right]_{i>j:(i,j)\in\mathcal{E}_{Y},l=1,\ldots,L_{ij}^{t}}\in\mathbb{R}^{L_{\mathrm{total}}}.$
The key step in establishing Theorems 2 and 3 is the following refined characterization of the estimation error $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$.
Lemma 4.
Instate the assumptions of Theorem 1. With probability at least $1-O(n^{-10})$, the estimator $\widehat{\bm{\theta}}$ given by the Algorithm (1) can be written as
$\widehat{\bm{\theta}}-\bm{\theta}^{\star}=-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}+\bm{r},$ (15)
As the residual term $\bm{r}$ has small magnitude, we may analyze the properties of $\widehat{\bm{\theta}}-\bm{\theta}^{\star}$ by focusing on the leading term $-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$.
Proof outline of Theorem 2.
Using the approximation $\widehat{\bm{\theta}}-\bm{\theta}^{\star}\approx-\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$, we can turn our attention to the concentration of
$\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|^{2}\approx\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}=\widehat{\bm{\epsilon}}^{\top}\bm{B}^{\top}\bm{L}_{Lz}^{\dagger}\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}.$
This is a quadratic form of sub-Gaussian random variables. The Hanson-Wright inequality tells us that $\|\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}\|^{2}$ concentrates around its expectation—$\mathrm{Trace}(\bm{L}_{Lz}^{\dagger})$. The complete proof is deferred to Section D for Theorem 2.
Proof outline of Theorem 3.
With Lemma 4 in hand, we can expand $\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ to reach
$\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})$ $\approx-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}$
$=-\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}\underbrace{\left[Y_{ji}^{(l)}-\sigma(\theta_{i}^{\star}-\theta_{j}^{\star})\right](\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})}_{\eqqcolon\bm{u}_{ij}^{(l)}}.$
The main component of the last expression is a sum of bounded random variables $\bm{u}_{ij}^{(l)}$ that are independent conditional on $\mathcal{G}_{Y}$. This sum has zero mean and variance
$\sum_{i>j:(i,j)\in\mathcal{E}_{Y}}\sum_{l=1}^{L_{ij}}z_{ij}(\bm{L}_{Lz}^{\dagger})^{1/2}(\bm{e}_{i}-\bm{e}_{j})(\bm{e}_{i}-\bm{e}_{j})^{\top}(\bm{L}_{Lz}^{\dagger})^{1/2}=\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}.$
Therefore applying multivariate CLT, we obtain
$-(\bm{L}_{Lz}^{\dagger})^{1/2}\bm{B}\widehat{\bm{\epsilon}}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,\bm{I}_{m}-\frac{1}{m}\bm{1}_{m}\bm{1}_{m}^{\top}\right).$
See the complete proof in Section E.
4.4 Proof outline of Lemma 4
Due to the importance and usefulness of Lemma 4, we provide a sketch of its proof and leave the full one to Section C.
The proof is inspired by the proof of Theorem 1 in [Che23], which analyzes MLE via the trajectory of the preconditioned gradient descent (PGD) dynamic starting from ground truth. More precisely, letting $\bm{\theta}^{0}=\bm{\theta}^{\star}$, we consider the PGD iterates defined by
@@ 4.4 Proof outline of Lemma 4
where $\eta>0$ is the step size of PGD. [Che23] shows that this dynamic converges to $\widehat{\bm{\theta}}$. We proceed one step further by establishing precise distributional characterization of $\hat{\bm{\theta}}$ via analyzing PGD. With Taylor expansion, the gradient can be decomposed into
$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
@@
We establish Lemma 4 by solving this recursive relation. More specifically, as $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$ does not depend on $t$ and $\|\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\|_{\infty}$ can be controlled for each step $t$, taking $t\rightarrow\infty$, we see that
@@
5 Experiment
In this section, we demonstrate the empirical performance of $\mathsf{RP\text{-}MLE}$ and its variant $\mathsf{MRP\text{-}MLE}$ using both simulated and real data.
5.1 Simulations
We begin by using simulated data to validate our theoretical results and compare our estimators with existing ones for the Rasch model.
The data generating process follows the model specified in Section 2.1. Unless specified otherwise, in each trial, the user and item parameters are randomly drawn from
5.1.1 $\ell_{\infty}$ estimation error
Here, we investigate the $\ell_{\infty}$ estimation error. First, we validate the theoretical result in $\ell_{\infty}$ estimation error of $\mathsf{RP\text{-}MLE}$. Second, we show how much advantage the $\mathsf{MRP\text{-}MLE}$ brings through multiple runs of data splitting. Lastly, we compare our methods with existing comparison-based algorithms, including the case where $\kappa_{1},\kappa_{2}$ are large.
Validating the theoretical result.
Theorem 1 tells us that the $\ell_{\infty}$ error scales as $1/\sqrt{np}$. As shown in Figure 1, $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ is very close to being linear w.r.t. both $1/\sqrt{n}$ and $1/\sqrt{p}$. This is consistent with our theoretical predictions.
(a) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ with v.s. $1/\sqrt{n}$. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $10000$ to $40000$. Each point represents the average of 1000 trials.
(b) $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $1/\sqrt{p}$. The parameter is chosen to be $m=50,n=10000$ and $p$ varies from $1/9$ to $1$. Each point represents the average of 1000 trials.
Figure 1: Estimation error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ of $\mathsf{RP\text{-}MLE}$ with varying $n$ and $p$.
Multiple runs in $\mathsf{MRP\text{-}MLE}$.
As we have discussed after Theorem 1, the random data splitting could incur a small loss of information. We have provided a remedy $\mathsf{MRP\text{-}MLE}$ (Algorithm 2) by averaging over multiple runs with independent data splitting. In Figure 2, we show the empirical evidence of its effectiveness. For each trial, we run 50 independent data splitting and obtain the 50 estimates $\widehat{\bm{\theta}}^{(i)}$. We then report the $\ell_{\infty}$ error of the running average, i.e., $\|\frac{1}{k}\sum_{i=1}^{k}\widehat{\bm{\theta}}^{(i)}-\bm{\theta}^{\star}\|$ for $k=1,\ldots,50$. As a baseline for comparison, we also run pseudo MLE on the same datasets, which uses all possible pairs of comparison. We can see that by averaging over more runs of data splittings, $\mathsf{MRP\text{-}MLE}$ gives a better estimation. In particular, one can see that with sufficient number of independent data splittings, the average estimation error matches and even beats pseudo MLE.
Figure 2: The $\ell_{\infty}$ estimation error of $\mathsf{MRP\text{-}MLE}$ with different numbers of data splitting. The dashed line is the performance of Pseudo MLE. The parameters are chosen to be $m=50,p=0.1,n=10000$. Each point is averaged over 1000 trials.
Comparison with existing estimator.
We then compare our algorithms with two other comparison-based algorithms: pseudo MLE (PMLE) and the spectral method in [NZ23]. We plot the $\ell_{\infty}$ error $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ against $n$ in Figure 4. We can see that all four algorithms follow a similar trend. The performance of $\mathsf{MRP\text{-}MLE}$ is on par with PMLE and better than the spectral method. Our proposed algorithm not only enjoys an improved theoretical guarantee but is also competitive in practice.
Performance with large $\kappa_{1},\kappa_{2}$.
While we assume $\kappa=\max\{\kappa_{1},\kappa_{2}\}=O(1)$ is most of this article, the scenario with large $\kappa$ can be practically relevant. Here we compare the $\ell_{\infty}$ error of different methods when the condition number is large. Fixing $\kappa$, we draw the user and item parameters with
and compute $\bm{\zeta}^{\star}$ and $\bm{\theta}^{\star}$ by shifting $\widetilde{\bm{\zeta}}^{\star}$ and $\widetilde{\bm{\theta}}^{\star}$ to zero mean. Figure 4 shows the performance of different estimators as $\kappa$ varies. The MLE-based approaches including $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ achieve better $\ell_{\infty}$ error than the spectral method when $\kappa$ is large.
Figure 3: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $n$ using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=50,p=0.1$ and $n$ varies from $5000$ to $20000$. The result is averaged over 1000 trials.
Figure 4: $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|_{\infty}$ v.s. $\log(\kappa)$ using Spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=50,p=0.1,n=20000$ and $\kappa$ varies from $1$ to $e^{10}$. The result is averaged over 1000 trials.
5.1.2 Top-$K$ recovery
In this part, we investigate the performance of different algorithms in top-$K$ recovery. We set $\theta_{i}^{\star}=(1-K/m)\Delta_{K}$ for $i\leq K$ and $\theta_{i}^{\star}=(-K/m)\Delta_{K}$ otherwise. For any estimator $\widehat{\bm{\theta}}$, we define top-$K$ recovery rate to be
$\frac{1}{K}\left|\left\{i\leq K:i\in A_{K}\right\}\right|,$
where $A_{K}$ is an arbitrary $K$-element set such that $\widehat{\theta}_{i}\geq\widehat{\theta}_{j}$ for any $i\in A_{K},j\notin A_{K}$. We compare the top-$K$ recovery rate of PMLE and the spectral method in [NZ23] with $\mathsf{RP\text{-}MLE}$ and $\mathsf{MRP\text{-}MLE}$ in Figure 6. The recovery rate of PMLE, spectral method and $\mathsf{MRP\text{-}MLE}$ is similar, indicating again that our algorithm performs well in practice.
Figure 5: Top-$K$ recovery rate using spectral method, PMLE, $\mathsf{RP\text{-}MLE}$, and $\mathsf{MRP\text{-}MLE}$ using 20 data splittings. The parameter is chosen to be $m=10000,m=50,p=0.1,K=5$ and $\Delta_{K}$ varies from 0.1 to 0.7. The result is averaged over 1000 trials.
5.1.3 Refined $\ell_{2}$ estimation error
In this part of the section, we show the empirical evidence of Theorem 2. In Theorem 2 we have shown that the $\ell_{2}$ error concentrate around $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ for $\widetilde{z}\in\left\{\widehat{z},z\right\}$. In each trial we compute the following quantity
In other words, $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ concentrate tightly around $\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})$. We can see that the deviation is very small between $\widetilde{z}=\widehat{z}$ and $\widetilde{z}=z$. In both cases, the relative deviation of $\|\widehat{\bm{\theta}}-\bm{\theta}^{\star}\|$ from $\sqrt{\mathrm{Trace}(\bm{L}_{L\tilde{z}}^{\dagger})}$ decreases as $n$ and $m$ increase as expected.
5.1.4 Inference
Lastly, we validate our inferential results in Section 3.3.
Asymptotic normality.
We first check the asymptotic normality presented in Theorem 3 and Corollary 1. For simplicity, we focus on the first coordinate. Theorem 3 and Corollary 1 have claimed that
@@ 5.1.4 Inference
$\left[\bm{L}_{Lz}^{1/2}(\widehat{\bm{\theta}}-\bm{\theta}^{\star})\right]_{1}\overset{\mathrm{d}}{\rightarrow}\mathcal{N}\left(0,1-\frac{1}{m}\right)$
and

### v2

@@ Appendix A Analysis
In this section, we present the main steps to obtain theoretical results in the previous section. Section A.1 provides a complete argument on the reduction to the BTL model we mentioned in Section 2.2. Section A.2 provides the analysis of the $\ell_{\infty}$ error, Section A.3 provides the analysis of the non-asymptotic expansion, and Section A.4 sketches the proof of the asymptotic normality of $\mathrm{MRP\text{-}MLE}$ and $\mathrm{WP\text{-}MLE}$.
A.1 Reduction to Bradley-Terry-Luce model
@@
A key component in $\mathrm{RP\text{-}MLE}$ is the random pairing in Steps 1 and 2 of Algorithm 1. It compiles the user-item responses $\bm{X}$ to item-item comparisons $\bm{Y}$. In this section, we make a detailed argument that $\bm{Y}$ follows the Bradley-Terry-Luce model with a non-uniform sampling scheme.
@@
Recall that $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}$ and $Y_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}<X_{tj}\}$. The following fact provides the distribution of $Y_{ij}^{t}$ conditional on $L_{ij}^{t}=1$. We defer its proof to Section E.1.
@@
    Randomly split the $m_{t}$ problems taken by user $t$ into $\lfloor m_{t}/2\rfloor$ pairs of problems. (Step 1(a) of Algorithm 1)
@@
    For all $(i,j,t)$, items $i$ and $j$ are compared by user $t$ if $L_{ij}^{t}\coloneqq\mathds{1}\{X_{ti}\neq X_{tj}\}=1$.
@@
In addition, we would like to comment on another popular method conditional MLE, which can also be viewed as a reduction to a item-only model. The CMLE maximizes the likelihood conditioned on total number of positive responses. It can be computed that
$\mathbb{P}\left[X_{ti_{1}},X_{ti_{2}},\cdots,X_{ti_{m_{t}}}\mid\sum_{l=1}^{m_{t}}X_{ti_{l}}=k\right]=\frac{\prod_{l=1}^{m_{t}}e^{\theta_{i_{l}}^{\star}}\mathds{1}\{X_{ti_{l}}=1\}}{\sum_{\bm{\alpha}\in\{0,1\}^{m_{t}}}\prod_{l=1}^{m_{t}}e^{\theta_{i_{l}}^{\star}}\mathds{1}\{\alpha_{l}=1\}}.$
It is easy to see that the conditional probability is not dependent on the user parameters $\bm{\zeta}^{\star}$. However, the model CMLE reduces to is less studied than the BTL model, especially in the setting of sparse observations.
A.2 Analysis for entrywise error bound
We have seen that analyzing $\mathrm{RP\text{-}MLE}$ under the Rasch model can be reduced to analyzing the MLE under the BTL model. This reduction allows us to invoke the result in the recent work [47] established for MLE in the BTL model with a general comparison graph.
@@ A.2 Analysis for entrywise error bound
The following lemma adapts Theorem 3 of the recent work [47] to our setting.
Lemma 1 (Theorem 3 in [47]).
@@
$\frac{np}{4\kappa_{2}}\leq\lambda_{m-1}(\bm{L}_{L})\leq\lambda_{1}(\bm{L}_{L})\leq 3np,$ (18)
@@
A.2.1 Proof of Theorem 1
Now we are ready to prove Theorem 1. We focus on analyzing $\mathrm{RP\text{-}MLE}$, as the analysis of $\mathrm{MRP\text{-}MLE}$ follows immediately from the union bound of the different data splitting and the triangular inequality:
@@ A.2.1 Proof of Theorem 1
A.3 Analysis for non-asymptotic expansion
To make the main text concise, we provide a sketch of the proof of Theorem 2 and leave the full one to Section C.1.
The proof is inspired by the proof of Theorem 1 in [5], which analyzes MLE via the trajectory of the preconditioned gradient descent (PGD) dynamic starting from ground truth. More precisely, letting $\bm{\theta}^{0}=\bm{\theta}^{\star}$, we consider the PGD iterates defined by
@@ A.3 Analysis for non-asymptotic expansion
where $\eta>0$ is the step size of PGD. [5] shows that this dynamic converges to $\widehat{\bm{\theta}}$. We proceed one step further by establishing precise distributional characterization of $\widehat{\bm{\theta}}$ via analyzing PGD. With Taylor expansion, the gradient can be decomposed into
$\nabla\mathcal{L}(\bm{\theta}^{t})=\bm{L}_{Lz}(\bm{\theta}^{t}-\bm{\theta}^{\star})-\bm{B}\widehat{\bm{\epsilon}}+\bm{r}^{t},$
@@
We establish Theorem 2 by solving this recursive relation. More specifically, as $\bm{L}_{Lz}^{\dagger}\bm{B}\widehat{\bm{\epsilon}}$ does not depend on $t$ and $\|\bm{L}_{Lz}^{\dagger}\bm{r}^{t}\|_{\infty}$ can be controlled for each step $t$, taking $t\rightarrow\infty$, we see that
@@
A.4 Analysis for asymptotic normality
The analysis of the asymptotic normality when $m,p$ are fixed is standard for maximum likelihood estimators. Here we illustrate the idea on $\mathrm{RP\text{-}MLE}$ for one random splitting. By mean value theorem
@@ A.4 Analysis for asymptotic normality
$\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star})$ $=\sum_{t=1}^{n}\nabla\mathcal{L}_{k}^{(t)}(\widehat{\bm{\theta}}_{k}^{(n)})+\left[\int_{\tau=0}^{1}\sum_{t=1}^{n}\nabla^{2}\mathcal{L}_{k}^{(t)}(\bm{\theta}^{\star}+\tau(\widehat{\bm{\theta}}_{k}^{(n)}-\bm{\theta}^{\star}))\right]\mathrm{d}\tau(\bm{\theta}^{\star}-\widehat{\bm{\theta}}_{k}^{(n)})$